# Score Change of Variables

### v1

@@ A.4 Derivation of Weighted Loss Function
$\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t)t)^{\top}\right)\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)}.$ (31)
@@
Appendix B Derivation of the Generalized Sliced Score Matching Objective
@@
B.1 Score Matching in Transformed Space
@@ B.2 Normalization Considerations
B.3 Eliminating the Dependence on $\nabla_{\mathbf{x}}\log p(\mathbf{x})$
@@
Appendix C Quadratic GSSM Details
@@
C.1 Quadratic GSSM Loss
@@
C.2 Variance Reduction for Quadratic GSSM
@@
C.2.1 Expectation of the Quadratic Term
@@
C.2.2 Expectation of the Hessian Term
@@
C.2.3 Expectation of the Trace Term
@@
C.2.4 Final Variance-Reduced Loss
@@
C.3 Sampling from $\mathbf{A}\mathbf{x}$
@@
C.4 Additional Experimental Results
@@
Appendix D Deep Kernel Exponential Families with Generalized Sliced Score Matching
@@
D.1 Deep Kernel Exponential Family Models
@@
D.2 Quadratic Form of the GSSM Loss
@@
D.3 Quadratic Form of GSSM Loss
@@ D.4 Implementation Details
Appendix E Detailed Derivations for Chess Diffusion
E.1 SDE Parameters and Transformation
@@
E.2 Transformed Drift and Diffusion Coefficients
The transformed drift $\bar{\mathbf{f}}(\mathbf{y},t)$ and diffusion $\bar{\mathbf{G}}(\mathbf{y},t)$ are:
$\bar{f}_{i}(\mathbf{y},t)$ $=y_{i}\left[k_{i}-\sum_{j=1}^{12}y_{j}k_{j}\right],$ (95)
$\bar{G}_{ij}(\mathbf{y},t)$ =$=\begin{cases}g(t)y_{i}(1-y_{i}),&\text{if }i=j,\\ -g(t)y_{i}y_{j},&\text{if }i\neq j,\end{cases}$ (96)
where $k_{i}=f_{i}(\mathbf{x},t)+\frac{1}{2}g^{2}(t)(1-2y_{i})$.
E.3 Score Function Transformation
The score function $\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})$ in terms of $\mathbf{x}$ is:
$(\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y}))_{i}$ $=\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\ \mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right)$ (97)
$=\frac{1}{y_{13}}\sum_{j=1}^{12}\frac{\partial}{\partial x_{j}}\log p_{t}(\mathbf{x})+\frac{1}{y_{i}}\frac{\partial}{\partial x_{i}}\log p_{t}(\mathbf{x})+\frac{y_{i}-y_{13}}{y_{i}y_{13}},$ (98)
where $y_{13}=1-\sum_{j=1}^{12}y_{j}$.
E.4 Divergence Term
The divergence term $\nabla_{\mathbf{y}}\cdot\left(\bar{\mathbf{G}}(\mathbf{y},t)\bar{\mathbf{G}}(\mathbf{y},t)^{\top}\right)$ for the $i$-th component is:
$(\nabla_{\mathbf{y}}\cdot\left(\bar{\mathbf{G}}(\mathbf{y},t)\bar{\mathbf{G}}(\mathbf{y},t)^{\top}\right))_{i}$ $=g(t)^{2}y_{i}\left(2-2S_{y}+15S_{y^{2}}-15y_{i}\right),$ (99)
where $S_{y}=\sum_{j=1}^{12}y_{j}$ and $S_{y^{2}}=\sum_{j=1}^{12}y_{j}^{2}$.
E.5 Training Details

### v2

@@ B.5 Derivation of Weighted Loss Function
$\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right)\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)}.$ (56)
@@
Appendix C Derivation of the Generalized Sliced Score Matching Objective
@@
C.1 Score Matching in Transformed Space
@@ C.2 Normalization Considerations
C.3 Eliminating the Dependence on $\nabla_{\mathbf{x}}\log p(\mathbf{x})$
@@
Appendix D Quadratic GSSM Details
@@
D.1 Quadratic GSSM Loss
@@
D.2 Variance Reduction for Quadratic GSSM
@@
D.2.1 Expectation of the Quadratic Term
@@
D.2.2 Expectation of the Hessian Term
@@
D.2.3 Expectation of the Trace Term
@@
D.2.4 Final Variance-Reduced Loss
@@
D.3 Sampling from $\mathbf{A}\mathbf{x}$
@@
D.4 Additional Experimental Results
@@
Appendix E Deep Kernel Exponential Families with Generalized Sliced Score Matching
@@
E.1 Deep Kernel Exponential Family Models
@@
E.2 Quadratic Form of the GSSM Loss
@@
E.3 Quadratic Form of GSSM Loss
@@ E.4 Implementation Details
Appendix F Chess Diffusion Details
F.1 SDE Parameters and Transformation
@@
F.2 Transformed Drift and Diffusion Coefficients
The transformed reverse drift $\hat{\mathbf{f}}(\mathbf{y},t)$ and diffusion $\tilde{\mathbf{G}}(\mathbf{y},t)$ are:
$\hat{f}_{i}(\mathbf{y},t)$ $=y_{i}\left[k_{i}-\sum_{j=1}^{12}y_{j}k_{j}\right],$ (120)
$\tilde{G}_{ij}(\mathbf{y},t)$ =$=\begin{cases}g(t)y_{i}(1-y_{i}),&\text{if }i=j,\\ -g(t)y_{i}y_{j},&\text{if }i\neq j,\end{cases}$ (121)
where $k_{i}=f_{i}(\phi^{-1}(\mathbf{y}),t)+\frac{1}{2}g^{2}(t)(2y_{i}-1)$.
F.3 Training Details